OpenStudy (lalaland_lauren):

If (6y − 11)(6y + 11) = ay2 − b, what is the value of a?

jimthompson5910 (jim_thompson5910):

hints: difference of square rule (x-y)*(x+y) = x^2 - y^2 (3z+2x)*(3z-2x) = (3z)^2 - (2x)^2 = 9z^2 - 4x^2 (2y+5)*(2y-5) = (4y)^2 - (5)^2 = 16y^2 - 25 none of these are the answer. They're just examples

OpenStudy (lalaland_lauren):

That didn't really help me :( can I just get the answer?

jimthompson5910 (jim_thompson5910):

Difference of squares rule \[\Large (\color{red}{x} - \color{blue}{y})(\color{red}{x} + \color{blue}{y}) = (\color{red}{x})^2 - (\color{blue}{y})^2\] \[\Large (\color{red}{6y} - \color{blue}{11})(\color{red}{6y} + \color{blue}{11}) = (\color{red}{6y})^2 - (\color{blue}{11})^2 = ???\] I'll let you do that last part

OpenStudy (lalaland_lauren):

Would it be 36y^2- 121?

jimthompson5910 (jim_thompson5910):

yes it would be

OpenStudy (lalaland_lauren):

so a would be 36

jimthompson5910 (jim_thompson5910):

correct